# Parkinson’s Disease Classification Using Contrastive Graph Cross-View Learning with Multimodal Fusion of SPECT Images and Clinical Features

### v3

@@ Abstract
Parkinson’s Disease (PD) affects millions globally, impacting movement. Prior research utilized deep learning for PD prediction, primarily focusing on medical images, neglecting the data’s underlying manifold structure. This work proposes a multimodal approach encompassing both image and non-image features, leveraging contrastive cross-view graph fusion for PD classification. We introduce a novel multimodal co-attention module, integrating embeddings from separate graph views derived from low-dimensional representations of images and clinical features. This enables more robust and structured feature extraction for improved multi-view data analysis. Additionally, a simplified contrastive loss-based fusion method is devised to enhance cross-view fusion learning. Our graph-view multimodal approach achieves an accuracy of 91% and an area under the receiver operating characteristic curve (AUC) of 92.8% in five-fold cross-validation. It also demonstrates superior predictive capabilities on non-image data compared to solely machine learning-based methods.
@@ 4.2 Ablation Study
In our research, we began by extracting image features using a CNN model and non-imaging DaTQUANT variables, followed by constructing two cross-views of the graph representation. We subsequently concatenated these different modalities to improve the predictive capability after model fusion. Table 2 showcases our experimental results from utilizing two methods (GCN and GAT) for learning graph structures and generating embeddings. The results revealed that, when a cross-view approach was employed, the GAT method achieved a macro average accuracy rate of 91%, along with F1, sensitivity, and precision scores of 92% in normal and abnormal classes. Additionally, our method can achieve an 5-fold cross-validated AUC of 92.8% as shown in (Fig. 2 (B)).

### v4

@@ Abstract
Parkinson’s Disease (PD) affects millions globally, impacting movement. Prior research utilized deep learning for PD prediction, primarily focusing on medical images, neglecting the data’s underlying manifold structure. This work proposes a multimodal approach encompassing both image and non-image features, leveraging contrastive cross-view graph fusion for PD classification. We introduce a novel multimodal co-attention module, integrating embeddings from separate graph views derived from low-dimensional representations of images and clinical features. This enables more robust and structured feature extraction for improved multi-view data analysis. Additionally, a simplified contrastive loss-based fusion method is devised to enhance cross-view fusion learning. Our graph-view multimodal approach achieves an accuracy of 0.91 and an area under the receiver operating characteristic curve (AUC) of 0.93 in five-fold cross-validation. It also demonstrates superior predictive capabilities on non-image data compared to solely machine learning-based methods.
@@ 4.2 Ablation Study
In our research, we began by extracting image features using a CNN model and non-imaging DaTQUANT variables, followed by constructing two cross-views of the graph representation. We subsequently concatenated these different modalities to improve the predictive capability after model fusion. Table 2 showcases our experimental results from utilizing two methods (GCN and GAT) for learning graph structures and generating embeddings. The results revealed that, when a cross-view approach was employed, the GAT method achieved a macro average accuracy rate of 0.91, along with F1, sensitivity, and precision scores of 0.92 in normal and abnormal classes. Additionally, our method can achieve an 5-fold cross-validated AUC of 0.93 as shown in (Fig. 2 (B)).